# Symmetry Breaking and Mie-tronic Supermodes in Nonlocal Metasurfaces

### v2

@@ III.2 $Q$ Enhancement via Symmetry Breaking
Bloch waves have long served as a powerful framework for describing electron behavior in crystals, underpinning much of the success of the modern electronics industry. However, it has been recognized since shortly after Bloch’s original formulation that, while Bloch waves offer an intuitive picture of electron transport, they neglect the atomic-scale structure of matter and therefore cannot describe electronic localization. This limitation led to the development of Wannier functions [Wannier1962dynamics], which now form the foundation of electronic-structure calculations for finite systems [Marrazzo2024wannier]. Although optical analogues of Wannier functions have been proposed for modeling light localization, the exponential confinement intrinsic to electronic Wannier functions is not directly applicable to electromagnetic waves. As shown in the accompanying Letter [Hoang2025unconventional], multipole functions provide a more natural and physically rigorous basis for investigating light localization within the Mie-tronic framework.
@@ IV Diffractive Nonlocal Metasurfaces
IV.1 Mie-Resonant Origin of Nonlocal Diffractive Bands
@@
In coupled arrays, both dipole and quadrupole modes split into bonding and antibonding branches. Within diffraction theory, only the antibonding modes couple efficiently to the radiation continua, giving rise to observable diffractive bands. Because dipole and quadrupole resonances occur at relatively low frequencies, their associated diffraction bands are spectrally well separated and tunable through geometry. Figures 6(b) and (c) illustrate this effect: reshaping the spherical unit cell into a hole-type geometry transforms the MD-based metasurface from reflection-dominated to transmission-dominated operation. The near-field maps obtained from RCWA (insets) confirm that these diffractive bands originate from the MD resonance.
@@
More fundamentally, the symmetry mismatch arises from destructive interference of multipolar displacement currents within the symmetric unit cell, which cancels the radiation of the MD supermode along the surface normal [Hoang2025collective]. This interference picture also explains how symmetry breaking restores coupling between P-polarized plane waves and the MD supermode, as demonstrated in the following subsection.
IV.2 Symmetry Breaking Enables Polarization Conversion
@@
V Discussion and Conclusion
Metasurfaces have become a cornerstone of modern photonics, offering miniaturized, chip-scale platforms for advanced optical functionalities. Yet, pushing their performance beyond current limits requires explicit consideration of nonlocal interactions among unit cells. Within the Mie-tronics framework, we have shown that symmetry breaking enables precise control over the coupling between metasurface supermodes and external excitations, leading to polarization conversion and enhanced light confinement in finite metastructures. These mechanisms lay the foundation for emerging functionalities such as analog optical computing and on-chip information processing.
@@ Acknowledgements.
This research is supported by the National Research Foundation, Singapore, and A*STAR under its Frontier Competitive Research Programme (NRF-F-CRP-2024-0009), the Australian Research Council (Grant No. DP210101292), and International Technology Center Indo-Pacific (ITC IPAC) via Army Research Office (Contract FA520923C0023). D.L. acknowledges a support from the Ministry of Education of Singapore under its SUTD Kickstarter Initiative (Grant No. SKI 20210501). J. Ji acknowledges the funding received from the PhotonDelta National Growth Fund programme.

### v3

@@ III.2 $Q$-enhancement via symmetry breaking
Bloch waves have long served as a powerful framework for describing electron behavior in crystals, underpinning much of the success of the modern electronics industry. However, it has been recognized since shortly after Bloch’s original formulation that, while Bloch waves offer an intuitive picture of electron transport, they neglect the atomic-scale structure of matter and therefore cannot describe electronic localization. This limitation led to the development of Wannier functions [Wannier1962dynamics], which now form the foundation of electronic-structure calculations for finite systems [Marrazzo2024wannier]. Although optical analogues of Wannier functions have been proposed for modeling light localization, the exponential confinement intrinsic to electronic Wannier functions is not directly applicable to electromagnetic waves. As shown in the accompanying Letter [Hoang2025unconventional], multipole functions provide a more natural and physically rigorous basis for investigating light localization within the Mie-tronics framework.
@@ IV Diffractive Nonlocal Metasurfaces
IV.1 Mie-resonance origin of nonlocal diffractive bands
@@
In coupled arrays, both dipole and quadrupole modes split into bonding and anti-bonding branches. Within diffraction theory, only the anti-bonding modes couple efficiently to the radiation continua, giving rise to observable diffractive bands. Because dipole and quadrupole resonances occur at relatively low frequencies, their associated diffraction bands are spectrally well separated and tunable through geometry. Figures 6(b) and (c) illustrate this effect: reshaping the spherical unit cell into a hole-type geometry transforms the MD-based metasurface from reflection-dominated to transmission-dominated operation. The near-field maps obtained from RCWA (insets) confirm that these diffractive bands originate from the MD resonance.
@@
More fundamentally, the symmetry mismatch arises from destructive interference of multipolar displacement currents within the symmetric unit cell, which cancels the radiation of the MD supermodes along the surface normal [Hoang2025collective]. This interference picture also explains how symmetry breaking restores coupling between P-polarized plane waves and the MD supermodes, as demonstrated in the following subsection.
IV.2 Polarization conversion via symmetry breaking
@@
V Discussion and Conclusions
Metasurfaces have become a cornerstone of modern photonics, offering future miniaturized, chip-scale platforms for advanced optical functionalities. Yet, pushing their performance beyond current limits requires explicit consideration of nonlocal interactions among unit cells. Within the Mie-tronics framework, we have shown that symmetry breaking enables precise control over the coupling between metasurface supermodes and external excitations, leading to polarization conversion and enhanced light confinement in finite-size metastructures. These mechanisms lay the foundation for emerging functionalities such as analog optical computing and on-chip information processing.
@@ Acknowledgements.
This research is supported by the National Research Foundation, Singapore, and A*STAR under its Frontier Competitive Research Programme (NRF-F-CRP-2024-0009), the Australian Research Council (Grant No. DP210101292). D.L. acknowledges a support from the Ministry of Education of Singapore under its SUTD Kickstarter Initiative (Grant No. SKI 20210501). J. Ji acknowledges the funding received from the PhotonDelta National Growth Fund programme.